OpenStudy (anonymous):

solve 4x-4y=-16, x-2y=-12 by using substitution. Express your answer as an ordered pair.

OpenStudy (anonymous):

x=-12+2y..(1) sub to 4x-4y=-16 4(-12+2y)-4y=-16 -48+8y-16y=-16 -8y=32 y=-4 sub to (1) x=-12-8=-20